I have a word problem with a few parts. Can you help me?
artic glaciers
owing to global warming the artic ice cap has been breaking up. As a result, there has been an increase in the number of ice bergs floating in the artic ocean. The table gives ab ice berg count "i" for various years " t".
T 1970 1980 2000
i. 400 600 1000
a) make a scatrer plot if the data.
B) fund the slope -intercept form of a line that models the number of icebergs "i" in year "t". Interpret the slope of this line.
C) is the line you found impart b) an exact model for the data in the table?
D) if current trends continue, what will be the iceberg count in 2005?

Answer by stanbon(75887)

artic glaciers
owing to global warming the artic ice cap has been breaking up.
As a result, there has been an increase in the number of ice bergs floating in the artic ocean.
The table gives ab ice berg count "i" for various years " t".
T 1970 1980 2000
i. 400 600 1000
a) make a scatter plot of the data.
Can't show a scatter plot on this site.
------
Put the T values on the horizontal axis
Plot a point the appropriate height above each year.
----------------------------------------------------------
B) find the slope -intercept form of a line that models the number of icebergs "i" in year "t". Interpret the slope of this line.
---
slope = (1000-600)/(2000-1980) = 400/20 = 20/1
Interpretation: For each addtional year there are 20 more icebergs.
Equation: i(t) = 20t + 400 where t is the number of years after 1970
-------------------------
C) is the line you found impart
????
---------------------------------------
b) an exact model for the data in the table?
i(t) = 20t + 400
-------------------------------
D) if current trends continue, what will be the iceberg count in 2005?
2005 is 35 years after 1970
i(35) = 20*35 + 400 = 700+400 = 1100 icebergs
==========================================
Cheers,
Stan H.